if you randomly synthesized RNA strands in vitro using a 1:3 mixture of C:U nucleotides, and then translated those RNA strands in vitro, what proportion of the amino acids in the resulting polypeptide chains will be serine amino acids?
Answer #1

Given that the ratio of C is 1/4 and U is 3/4.

So the frequency of C in the mRNA will be 1/4 = 0.25

So the frequency of U in the mRNA will be 3/4 = 0.75

Possible codon and their frequency -

Frequency of UUU  =0.75*0.75*0.75 = 0.422

UUC - 0.75*0.75*0.25 = 0.140

UCU - 0.75*0.75*0.25  = 0.140

CUU - 0.75*0.75*0.25 =0.140

CCU - 0.75*0.25*0.25 =0.0468

CUC- 0.75*0.25*0.25 =  0.0468

UCC- 0.75*0.25*0.25  = 0.0468

CCC - 0.25*0.25*0.25 =0.015

Total frequency - 1

UUU = Codes for phenylalanine - Its frequency will be 0.422

UUC - Codes for phenylalanine - Its frequency will be 0.140

UCU -  Codes for Serine - Its frequency will be 0.140

CUU -  Codes for Leucine - Its frequency will be 0.140

CCU -  Codes for Proline - Its frequency will be 0.0468

CUC- Codes for Leucine - Its frequency will be 0.0468

UCC-Codes for serine- Its frequency will be 0.0468

CCC - Codes for proline its frequency will be 0.015.

Please write to me if you need more help.